OpenStudy (anonymous):

What is the equation of the line passing through (2, -2) and perpendicular to the line y = -x + 3 in slope-intercept form?

OpenStudy (anonymous):

y = -x + 3 y = mx +b you have x & y solve for b

OpenStudy (anonymous):

Uhhhhhmmm,how though?

OpenStudy (anonymous):

y = mx +b perpendicular slope = -1/m = -1/-1 = 1 y = 1x +b you have (2,-2) -2 = 1(2) +b now?

OpenStudy (anonymous):

y = x - 4?

OpenStudy (anonymous):

yup -2 = 1(2) +b -2 -2 b = -4 y = x - 4